OpenStudy (anonymous):

A highway makes an angle of 6 with the horizontal. This angle is maintained for a horizontal distance of 5 miles. To the nearest hundredth of a mile, how high does the highway rise in this 5-mile section? Show the steps you use to find the distance.

OpenStudy (anonymous):

tan(angle) = opposite over adjacent, => tan(6°) = x/8 so x = 8 * tan(6°) = 0.84 miles